OpenStudy (anonymous):

(1-cos(t))/(sin(t))=csc(t)-cot(t)

OpenStudy (anonymous):

This one is rather simple:\[ \frac{1-\cos(t)}{\sin(t)} = \frac{1}{\sin(t)}-\frac{\cos(t)}{\sin(t)} = \csc(t)-\cot(t) \]

OpenStudy (anonymous):

Oh you just split it apart..well damn..thanks lol